OpenStudy (narissa):

Which number could be multiplied by each side of the equation to produce the equivalent equation x=10 4/5x=8 A 5/4 B -5/4 C1/5 D 5 -1/6X=7 A -1 B -42 C-6/7 D 7 1/6 4x+6=38 A 8 B 3.5 C 144 D 11

OpenStudy (narissa):

@MyNameIsNemo

OpenStudy (mynameisnemo):

okay 1 sec

OpenStudy (narissa):

kk

OpenStudy (mynameisnemo):

1. is A because 8 times 5/4 is 10 which produces X = 10

OpenStudy (mynameisnemo):

wait wait wait

OpenStudy (narissa):

okay lol

OpenStudy (narissa):

i think 2 is B and 3 is A not sure about 1 though

OpenStudy (mynameisnemo):

I'm going to need a minute

OpenStudy (narissa):

alright take your time

OpenStudy (mynameisnemo):

Take the medal you gave me away :(

OpenStudy (mynameisnemo):

i don't know the answer, i'll tag someone who can help you better

OpenStudy (narissa):

alright thanks

OpenStudy (narissa):

@IrishBoy123

OpenStudy (narissa):

@wwhitlock

OpenStudy (narissa):

@texaschic101

OpenStudy (narissa):

@malcolmmcswain

OpenStudy (narissa):

@freckles

OpenStudy (narissa):

@robtobey

OpenStudy (narissa):

is it 42

OpenStudy (freckles):

.\[\frac{a}{b}x=c \text{ can be solved for } x \\ \text{ by multiplying both sides by } \frac{b}{a} \\ \frac{b}{a} \cdot \frac{a}{b} x=\frac{b}{a} \cdot c \\ x=\frac{b}{a} \cdot c \\ x=\frac{b \cdot c}{a}\]

OpenStudy (freckles):

Which number could be multiplied by each side of the equation to produce the equivalent equation x=10 4/5x=8 A 5/4 B -5/4 C1/5 D 5 -1/6X=7 A -1 B -42 C-6/7 D 7 1/6 --- so looking at the first question how can you isolate x there? what do you need to multiply on both sides?

OpenStudy (narissa):

multiply both sides by 8?

OpenStudy (freckles):

it you multiply both sides by 8 you get 32/5x=64

OpenStudy (freckles):

this does not isolate x.

OpenStudy (narissa):

4

OpenStudy (freckles):

did you look at the example I did second ago.. all i did was multiply both sides by the reciprocal of the number in front of x

OpenStudy (freckles):

4/5x=8 4/5 is the number in front of x the reciprocal of 4/5 is?

OpenStudy (narissa):

5/4 so the answer is A right

OpenStudy (freckles):

yes \[\frac{4}{5} x=8 \\ \text{ multiply both sides by } \frac{5}{4} \\ \frac{5}{4} \frac{4}{5} x= \frac{5}{4} 8 \\ \frac{5}{5} \frac{4}{4} x=\frac{5(8)}{4} \\ (1)(1)x=\frac{40}{4} \\ x=10\]

OpenStudy (narissa):

Thanks :))

OpenStudy (freckles):

the other two equations...it is impossible to get x=10 because x=10 isn't a solution to the other two equations so that doesn't make sense unless the directions are different for the other two ?

OpenStudy (narissa):

okay so for the other question i think 2 is b and 3 is a

OpenStudy (freckles):

So is the directions for the other two just to solve because as I said it is impossible to get x=10 for the other two.

OpenStudy (narissa):

yeah we just have to solve

OpenStudy (freckles):

well yes you are right -42/-6 is 7... let's look at solving the last one

OpenStudy (freckles):

to solve linear equations in the form ax+b=c for x.... I usually start off with subtract b on both sides ax=c-b then I divide both sides by a x=(c-b)/a

OpenStudy (freckles):

4x+6=38 so what is your first step here?

OpenStudy (narissa):

subtract 6 from both sides 4x=32 4 divide by 32 = 8 x=8

OpenStudy (freckles):

looks great and 8 is right since 4*8+6 is 32+6 which is 38

OpenStudy (narissa):

alright thanks for all the help:)

OpenStudy (freckles):

np